OpenStudy (omarbirjas):

WILL MEDAL A multiple choice test has 10 questions. Each question has four answer choices. a. What is the probability a student randomly guesses the answers and gets exactly six questions correct? b. Is getting exactly 10 questions correct the same probability as getting exactly zero correct? Explain. c. Describe the steps needed to calculate the probability of getting at least six questions correct if the student randomly guesses. You do not need to calculate this probability! It would be greatly appreciated if you could show work as to how to solve these, thanks!

OpenStudy (omarbirjas):

@dan815 @Preetha

OpenStudy (omarbirjas):

@iGreen @TheSmartOne @thomaster

OpenStudy (omarbirjas):

@jigglypuff314 @Jaynator495

OpenStudy (omarbirjas):

@phi PLEASE HELP....

OpenStudy (misty1212):

exactly six out of ten?

OpenStudy (misty1212):

\[\binom{10}{6}\times \left(\frac{1}{4}\right)^6\times \left(\frac{3}{4}\right)^4\]

OpenStudy (misty1212):

probability you guess right is \(\frac{1}{4}\) probability you guess wrong is \(\frac{3}{4}\) guess right 6 times, wrong 4 times and \(\binom{10}{6}\) ways to do it

OpenStudy (misty1212):

b. Is getting exactly 10 questions correct the same probability as getting exactly zero correct? Explain. hell no very unlikely to guess all ten right since you have one in four chance of getting any one right, three in four of getting any one wrong

OpenStudy (misty1212):

c. Describe the steps needed to calculate the probability of getting at least six questions correct if the student randomly guesses. You do not need to calculate this probability! at least 6 means 6, 7, 8, 9, or 10 you would have to compute each of these probabilities and add them up

OpenStudy (omarbirjas):

yea b is kinda obvious and a stupid question...idk why they have it...

OpenStudy (omarbirjas):

I gave you a cookie.... god bless you.